# Functional Liftings of Restricted Geometric Inequalities

### v1

@@ Abstract.
We investigate what we term “generalized sup-convolutions”. We show that functional inequalities that enjoy an interpretation as sup-convolution inequalities can be deduced from the special case of indicator functions corresponding to a geometric inequality. As consequences we derive a Borell-Brascamp Lieb inequality for the Gaussian Brunn-Minkowski inequality and give a functional analog of the log-Brunn-Minkowski conjecture. We also show that Barthe’s reverse Brascamp-Lieb inequality is equivalent to the geometric inequality it implies, and give a functional formulation of a conjecture of Schneider. Though we focus on Euclidean applications, our results are general and can be directly applied in more abstract settings, like groups or even topological measure spaces without algebraic structure, we instantiate this claim with a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
@@ 1. Introduction
In this article we investigate the equivalence of certain types of geometric inequalities and their functional analogs, each of which can be reviewed as reverse Hölder type inequalities. Our setting will be abstract and rather flexible, but let us start with an example: the well-known equivalence of the celebrated Brunn-Minkowski inequality (see [66]), a linchpin for modern convex geometry, and the Prékopa-Leindler inequality (see [59, 38]), it’s functional version possessing broad applications in various fields throughout mathematics.
@@
where $|\cdot|$ (or $|\cdot|_{n}$) is the $n$-dimensional Lebesgue volume, is valid. The Prékopa-Leindler inequality asserts that for any triple of non-negative measurable functions $f,g$ and $h$ such that
@@
The numerous extensions and the ramifications of the Brunn-Minkowski and Prekopa-Leindler inequality has had in various fields of mathematics are detailed in the survey of Gardner [27]. An exhaustive list of references is beyond the scope of this paper. We would like to highlight two very important and classical papers of Borell [10], and Brascamp and Lieb [15], where generalizations of both of these inequalities have appeared; such inequalities have become referred to as “Borell-Brascamp-Lieb” and “Brascamp-Lieb” inequalities. More recent examples of such works, as well as modern treatments, can be found in the books [3, 4, 66].
We now return to our original point of investigation: the equivalence of the inequalities (1) and (3).
Taking $f=\mathbbm{1}_{A}$ and $g=\mathbbm{1}_{B}$, in (3), yields
@@
held for $A,B$ origin-symmetric convex bodies in ${\mathbb{R}}^{n}$ (see Section 6). By Borell [10], an $n$-dimensional probability measure that satisfies (5) for all $A,B$ Borel sets is necessarily the uniform measure on a bounded convex set, thus such an inequality necessarily fails for $\gamma$ and general $A,B$. Nevertheless this question was answered in the affirmative by Eskenazis and Moschidis [24]. Our first result is a functional version of (5). It says that with the additional assumption that the relevant functions are even and unimodal [^1] We consider a function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ unimodal if its super-level sets are convex bodies (see Section 2), the Borell-Brascamp-Lieb inequality of the Lebesgue measure holds for the Gaussian.
@@
Let $\gamma$ be the standard Gaussian probability measure on ${\mathbb{R}}^{n}$, let $t\in[0,1]$, $\alpha\geq-\frac{1}{n}$, let $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be even, unimodal functions. Assume that $h\colon{\mathbb{R}}^{n}\to[0,\infty)$ is any Borel measurable function for which
@@
We mention that [1, Theorem 1.7] is a special case of Theorem 1.1 when $f,g$ are assumed to be $\log$-concave and $\alpha\geq 0,$ which in turn is the Gaussian special case of Theorem 3.3 to come. The $\alpha=0$ case corresponds to (2) and (3). The key to our approach is the study of properties held by what we will refer to as a sup-convolution [^2] Note that this a multiplicative version of a sup-convolution, when $f=e^{-V},g=e^{-W}$, $\sup_{(1-t)x+ty=z}f^{1-t}(x)g^{t}(y)=\exp\left\{-\inf_{(1-t)x+ty=z}\bigg{[}(1-t)V(x)+tW(y)\bigg{]}\right\}.$ The quantity $\inf_{(1-t)x+ty=z}\bigg{[}(1-t)V(x)+tW(y)\bigg{]}$ is referred to as the inf-convolution of $V$ and $W$. . For instance in the $\alpha=0$ case, given a pair of non-negative measurable functions $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$, we set
@@
A property of the $\square$ operation that proves pivotal, beyond some reasonably expected monotonocity and measurability (see Section 2 for a complete development), is a somewhat peculiar, restricted reverse triangle inequality. That is given two pairs of functions $(f_{0},g_{0})$ and $(f_{1},g_{1})$ such that $f_{1}(x)>0$ implies that $f_{0}(x)\geq f_{0}(y)$ for all $y\in\mathbb{R}^{n}$, and $g_{1}(x)>0$ implies $g_{0}(x)\geq g_{0}(y)$ for all $y\in\mathbb{R}^{n}$, then
@@
We will see that the same behavior arises in various circumstances, for instance an analogous approach will yield the following result for the geometric mean of convex bodies (see Section 5 for further details).
@@
Let $t\in[0,1]$. Then the Lebesgue measure satisfies
$|A^{1-t}B^{t}|\geq|A|^{1-t}|B|^{t}$
for all even convex bodies $A$ and $B$, if and only if $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ even and unimodal functions, such that
@@
$\int_{\mathbb{R}^{n}}h\geq\left(\int_{\mathbb{R}^{n}}f\right)^{1-t}\left(\int_{\mathbb{R}^{n}}g\right)^{t}.$
That is, Theorem 1.2 gives an equivalent functional version of the celebrated $\log$-Brunn-Minkowski conjecture of Böröczky, Lutwak, Yang, and Zhang [14]. In particular, Theorem 1.2 affirmatively answers a question posed by Andrea Colesanti “Is there a functional version of the log-Brunn-Minkowski inequality?” at the conference 61 Probability Encounters, in honour of Sergey Bobkov held at Paul Sabatier University, Toulouse, France in the summer of 2023 [17, Main Question].
Moreover, in the unconditional case [^3] A function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ is said to be unconditional if, for each $\varepsilon_{i}\in\{-1,1\}$ $i=1,\dots,n$, and $x=(x_{1},\dots,x_{n})\in{\mathbb{R}}^{n}$, $f(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})=f(x)$; a set is unconditional if its characteristic function is an unconditional function., where the log-Brunn-Minkowski conjecture has been verified (see [63]), we have the following result.
@@
Let $t\in[0,1]$. For $f=(f_{1},f_{2})$, with $f_{1},f_{2}\colon{\mathbb{R}}^{n}\to[0,\infty)$ even, unconditional, and unimodal functions, and suppose that $h$ satisfies
$h(z)\geq f_{1}(u)^{1-t}f_{2}(v)^{t},$
for $(u,v)$ such that $z\in\{f_{1}\geq f_{1}(u)\}^{1-t}\{f_{2}\geq f_{2}(v)\}^{t}$. Then the $\log$-Prékopa-Leindler inequality holds:
(7) $\int_{{\mathbb{R}}^{n}}h\geq\left(\int_{{\mathbb{R}}^{n}}f_{1}\right)^{1-t}\left(\int_{{\mathbb{R}}^{n}}f_{2}\right)^{t}.$
In fact, all of the above results can be considered corollaries of Theorem 2.3 below. Among other implications to come is the following result on the saturation of Barthe’s reverse Brascamp-Lieb inequality by the geometric inequality it implies.
Theorem 1.4.
Let $n,m$ be positive integers. Consider integers $1\leq n_{1},\dots,n_{m}<n$ and constants $c_{1},\dots,c_{m}>0$ such that $\sum_{i}c_{i}n_{i}=n$, and let $c=\sum_{i}c_{i}$, and for each $i$, set $t_{i}=\frac{c_{i}}{c}$. For each $i=1,\dots,m$, assume that $B_{i}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n_{i}}$ is a surjective linear map, and that $\cap_{i}\text{ker}(B_{i})=\{0\}$.
Barthe’s reverse Brascamp-Lieb inequality (see Inequality (10) below) is equivalent to the following geometric inequality: for any collection of Borel sets $A_{1},\dots,A_{m}$, with $A_{i}\subset{\mathbb{R}}^{n_{i}}$, we have
$\left|\sum_{i=1}^{m}c_{i}B_{i}^{*}A_{i}\right|\geq E\mathcal{M}_{0}^{(c)}(|A_{i}|_{n_{i}}).$
We refer the reader to Section 2, Equation (8) for the definition of $\mathcal{M}_{0}^{(c)}$ and to Section 4, Equation (11), for that of the constant $E$. and note that this saturation is anticipated by the recent work of [13] who have already shown the set theoretic case to be extremal in the special case of “geometric instances”.
Further we show the equivalence of another conjecture and a functional analog. More explicitly Schneider conjectured in [65] that among convex bodies of a fixed volume, the volume of the $m$-th order difference body $D^{m}(K)$ (see Conjecture 4.4) is minimized by ellipsoids when $n\geq 3$ and $m\geq 2$. We show in Section 4 that this is equivalent to the following functional inequality for $f:\mathbb{R}^{n}\to[0,\infty)$ unimodal,
$\int_{(\mathbb{R}^{m})^{n}}\left(\sup_{x\in\mathbb{R}^{n}}\prod_{i=0}^{m}f(x-z_{i})\right)^{\frac{1}{m+1}}dz\geq c(m,n)\left(\int_{\mathbb{R}^{n}}f^{\frac{1}{m}}\right)^{m},$
where $c(m,n)\coloneqq|D^{m}(B)|_{nm}|B|^{m}$ and $B$ is the $\ell_{2}$ unit ball. In the same article, Schneider proved that the volume of the $m$-th order difference body for $n=2$ and any $m$ is minimized by centrally symmetric convex bodies.
As mentioned in the abstract the techniques developed are not limited to Euclidean spaces. In fact, the ideas here are inspired by, and could be used to recover, the proof of a Prekopa-Leindler type inequality on the integers, by the second author and Marsiglietti [55, Theorem 3.8]. To further hint at broader application, we also give a Borell-Brascamp-Lieb inequality for nilpotent groups, generalizing Tao [67], which in turn generalized the case given for the Heisenberg group [56].
Theorem 1.5.
Let $G$ be a simply connected nilpotent Lie group of (topological) dimension ddwith a Haar measure $\mu$ and $\mathcal{M}_{p}^{(t)}(u,v)$ the $t$-weighted $p$-mean [^4] We refer to the next section and more precisely to (8) for the definition of $\mathcal{M}_{p}^{(t)}$. of $u$ and $v$. Then, for $\alpha\in[-\frac{1}{d},1]$ and $t\in(0,1)$, if
@@
Let us outline the remainder of the article. In Section 2 we introduce an abstract sup-convolution, and derive our main technical tool in Theorem 2.3. In Section 3 we derive the dimensional Brunn-Minkowski inequality for the Gaussian measure and generalizations thereof. In Section 4 we deal with the equivalences betweeen set theoretic and function theoretic formulations of Barthe’s reverse Brascamp-Lieb inequality and Schnieder’s conjecture. In Section 5 we give a functional version of the log-Brunn-Minkowski conjecture and related results for the $L_{p}$-Brunn-Minkowski theory. Finally in Section 6 we discuss further generalizations of the development in Section 2, limitations of our methods, and give a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
@@ 2. Generalized Sup-Convolutions
For $u=(u_{1},\dots,u_{n})$ with $u_{i}>0$, $t=(t_{1},\dots,t_{n})$ with $t_{i}>0$ and $p\in(-\infty,0)\cup(0,\infty)$ we define the “mean” ($\sum_{i}t_{i}$ is not necesarily set to be one)
@@
is a sup-convolution as well. This in a very particular case of the next result.
@@
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$, let
$\mathcal{M}_{\alpha}^{(t)}(u_{i})=\left(\sum_{i=1}^{n}t_{i}u_{i}^{\alpha}\right)^{\frac{1}{\alpha}}$
and $C(z)=\{x:\Phi(x)=z\}$ for a continuous function $\Phi:E=E_{1},\times\cdots\times E_{n}\to F$. The map defined by
@@
We proceed to verifying superaddivity. Assume that $f=(f_{1},\dots,f_{n})$ and $g=(g_{1},\dots,g_{n})$, with $f_{i},g_{i}$ being step functions, such that if $g(x_{i})>0$, then $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. Now if $\square g(z)=0$, then the result is trivial from monotonicity. Otherwise, $g(z)>0$, and there exists [^5] Since $g_{i}$ are step functions $x\mapsto\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))$ takes only finite many values. an $x\in C(z)$ such that $\square g(z)=\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))>0$. For such an $x$, necessarily $g_{i}(x_{i})>0$ so that $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. It follows that
@@
Then $\psi(P)=\Phi(A)$, so that $\Phi(A)$ is analytic, and universally measurable (see [26, Chapter 43]). It follows that $\square s$ is universally measurable as well. ∎
@@
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$, a triple $(\nu,\mu,\square)$ and $p=(p_{1},\dots,p_{n})\in[0,\infty)^{n}$ satisfies
$\nu(\square a)\geq\mathcal{M}_{\alpha}^{(t)}(\|a_{i}\|_{p_{i}})$
@@
$\nu_{*}(\square f)\geq\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}}),$
@@
We prepare for the proof with the next lemma.
@@
In what follows, for $p=(p_{1},\dots,p_{n})$ with $p_{i}>0$ we denote for brevity
$\|f\|_{p}\coloneqq\left(\|f_{1}\|_{p_{1}},\dots,\|f_{n}\|_{p_{n}}\right).$
@@
Let us first observe that it is enough to prove the inequality for $f_{i}$ step functions. Define
@@
$\nu_{*}\left(\square f\right)\geq\nu\left(\square f^{(m)}\right)\geq\mathcal{M}_{\alpha}^{(t)}\left(\|f^{(m)}\|_{p}\right),$
@@
$\mathcal{M}_{\alpha}^{(t)}(\|f\|_{p})=\lim_{m\to\infty}\mathcal{M}_{\alpha}^{(t)}(\|f^{(m)}\|_{p})\leq\nu(\square f).$
@@
Note that $f_{1}^{-}$ and $f_{1}^{+}$ individually take strictly fewer values than $f_{1}$. Define
$\lambda=\frac{\|f^{-}_{1}\|_{p_{1}}}{\|f_{1}\|_{p_{1}}},$
so that $1-\lambda=\frac{\|f^{+}_{1}\|_{p_{1}}}{\|f_{1}\|_{p_{1}}}$. Note that $\lambda\in(0,1)$. Let
$f_{i}^{-}=\min\{f_{i},t_{i}\}\ \hbox{ and }\ f_{i}^{+}=\max\{f_{i}-t_{i},0\}.$
where $t_{i}$ is implicitly and uniquely defined through the equation
$\frac{\|f_{i}^{-}\|_{p_{i}}}{\|f_{i}\|_{p_{i}}}=\lambda$
because $t\mapsto\mu_{i}(\min\{f_{i},t\})$ is continuous and strictly increasing on $(0,\|f_{i}\|_{\infty})$.

### v3

@@ Abstract.
We investigate what we term “generalized sup-convolutions”. We show that functional inequalities that enjoy an interpretation as sup-convolution inequalities can be deduced from the special case of indicator functions corresponding to a geometric inequality. As consequences we derive a Borell-Brascamp Lieb inequality for the Gaussian Brunn-Minkowski inequality and give a functional analog of the log-Brunn-Minkowski conjecture. Though we focus on Euclidean applications, our results are general and can be directly applied in more abstract settings, like groups or even topological measure spaces without algebraic structure, we instantiate this claim with a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
@@ 1. Introduction
In this article we investigate the equivalence of certain types of geometric inequalities and their functional analogs, each of which can be reviewed as reverse Hölder type inequalities. Our setting will be abstract and rather flexible, but let us start with an example: the well-known equivalence of the celebrated Brunn-Minkowski inequality (see [55]), a linchpin for modern convex geometry, and its functional version, the well-known Prékopa-Leindler inequality (see [50, 31]), possessing broad applications in various fields throughout mathematics. We now present these two inequalities.
@@
where $|\cdot|$ (or $|\cdot|_{n}$) is the $n$-dimensional Lebesgue volume, is valid. The Prékopa-Leindler inequality asserts that for any triple of non-negative measurable functions $f,g$ and $h$ on $\mathbb{R}^{n}$ such that
@@
The numerous extensions and the ramifications of the Brunn-Minkowski and Prékopa-Leindler inequality has had in various fields of mathematics are detailed in the survey of Gardner [24]. An exhaustive list of references is beyond the scope of this paper. We would like to highlight two very important and classical papers of Borell [7], and Brascamp and Lieb [11], where generalizations of both of these inequalities have appeared; such inequalities have become referred to as “Borell-Brascamp-Lieb” and “Brascamp-Lieb” inequalities. More recent examples of such works, as well as modern treatments, can be found in the books [3, 4, 55].
We now return to our original point of investigation: the equivalence of the inequalities (1) and (3). Taking $f=\mathbbm{1}_{A}$ and $g=\mathbbm{1}_{B}$, in (3), yields
@@
held for $A,B$ origin-symmetric convex bodies in ${\mathbb{R}}^{n}$ (see Section 5) and $\gamma$ the standard Gaussian measure on $\mathbb{R}^{n}$. By Borell [7], an $n$-dimensional probability measure that satisfies (5) for all $A,B$ Borel sets is necessarily the uniform measure on a bounded convex set, thus such an inequality necessarily fails for $\gamma$ and general $A,B$. Nevertheless this question was answered in the affirmative by Eskenazis and Moschidis [21].
More generally, it is conjectured that an improvement like (5), in the class of symmetric convex bodies, holds for any even log-concave measure $\mu$ on $\mathbb{R}^{n}$. The most general result in this direction is due to Cordero-Erausquin and Rotem [17], where it is proved that (5) holds for rotationally invariant measures $d\mu(x)=e^{-w(|x|)}dx,$ where $w:[0,\infty)\to(-\infty,\infty]$ is increasing with $\mathbb{R}\ni t\mapsto w(e^{t})$ convex. Note that this class goes beyond log-concavity as it contains, e.g., the Cauchy measure.
Our first result is a functional version of the result in [17]. We say that a function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ unimodal if its super-level sets $\{f\geq t\}$ are convex bodies (see Section 2).
@@
Let $\mu$ be the measure on $\mathbb{R}^{n}$, with density $e^{-w(|x|)}$, where $w:[0,\infty)\to(-\infty,\infty]$ is increasing and $\mathbb{R}\ni t\mapsto w(e^{t})$ is convex. Let $t\in(0,1)$, $\alpha\geq-\frac{1}{n}$. If $f,g:\mathbb{R}^{n}\to[0,\infty)$ are even unimodal and $h:\mathbb{R}^{n}\to[0,\infty)$ is Borel measurable satisfying
(6) $h((1-t)x+ty)\geq[(1-t)f(x)^{\alpha}+tg(y)^{\alpha}]^{\frac{1}{\alpha}},\qquad\qquad\forall x,y\in\mathbb{R}^{n},$
then
(7) $\int_{\mathbb{R}^{n}}h\ d\mu\geq\left((1-t)\left(\int_{\mathbb{R}^{n}}f\ d\mu\right)^{\beta}+t\left(\int_{\mathbb{R}^{n}}g\ d\mu\right)^{\beta}\right)^{\frac{1}{\beta}},$
where $\beta=\frac{\alpha}{1+n\alpha}$. Moreover, (7) is equivalent to the geometric inequality
(8) $\mu\left((1-t)K+tL\right)^{\frac{1}{n}}\geq(1-t)\mu(K)^{\frac{1}{n}}+t\mu(L)^{\frac{1}{n}},$
for all $K,L\subset{\mathbb{R}}^{n}$ symmetric convex bodies.
Very recently, Cordero-Erausquin and Eskenazis [16] found a functional analogue of (8) by following a novel geometric way to use the so-called $L^{2}$-method. In Section 3, we explain how Theorem 1.1 implies their main result, while answering [16, Question 19] in the radially symmetric case.
In another recent work [1] functional analogues of the said geometric inequalities for smaller class of measures were proven via completely different tools coming from information theory. Specializing Theorem 1.1 to the Gaussian measure we are able to generalize [1, Theorem 1.7] when $f,g$ are assumed to be $\log$-concave and $\alpha\geq 0.$
Corollary 1.2.
Let $\gamma$ be the standard Gaussian probability measure on ${\mathbb{R}}^{n}$, $t\in[0,1]$ and $\alpha\geq-\frac{1}{n}$. Let $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$ be even, unimodal functions. Assume that $h\colon{\mathbb{R}}^{n}\to[0,\infty)$ is any Borel measurable function for which
@@
We also recover [1, Theorem 1.8], a functional analogue of a dimensional Brunn–Minkowski type inequality for star bodies, see Section 3.
Our approach [^1] We direct the reader to [15] for other approaches. can be traced back at least to the 1956 proof of the Brunn-Minkowski inequality by Hadwidger and Ohmann. It was their “slicing and balancing” of sets that we adapted to functions. In particular to what we will refer to as sup-convolutions [^2] Note that this a multiplicative version of what is usually called a sup-convolution, when $f=e^{-V},g=e^{-W}$, $\sup_{(1-t)x+ty=z}f^{1-t}(x)g^{t}(y)=\exp\left\{-\inf_{(1-t)x+ty=z}\bigg[(1-t)V(x)+tW(y)\bigg]\right\}.$ The quantity $\inf_{(1-t)x+ty=z}\bigg[(1-t)V(x)+tW(y)\bigg]$ is referred to as the inf-convolution of $V$ and $W$. . For instance in the $\alpha=0$ case, given a pair of non-negative measurable functions $f,g\colon{\mathbb{R}}^{n}\to[0,\infty)$, we set
@@
A property of the $\square$ operation that proves pivotal, beyond some reasonably expected monotonicity and measurability (see Section 2 for a complete development), is a somewhat peculiar, restricted reverse triangle inequality. That is given two pairs of functions $(f_{0},g_{0})$ and $(f_{1},g_{1})$ such that $f_{1}(x)>0$ implies that $f_{0}(x)\geq f_{0}(y)$ for all $y\in\mathbb{R}^{n}$, and $g_{1}(x)>0$ implies $g_{0}(x)\geq g_{0}(y)$ for all $y\in\mathbb{R}^{n}$, then
@@
A famous conjecture in the geometric of convex bodies is the $\log$-Brunn-Minkowski conjecture put forth by Böröczky, Lutwak, Yang and Zhang in [10]: given a pair of origin symmetric convex body $K,L\subset{\mathbb{R}}^{n}$ and $t\in(0,1)$, is it necessarily true that
(10) $|K^{1-t}L^{t}|\geq|K|^{1-t}|L|^{t}?$
Here $K^{1-t}L^{t}$ is the $L_{0}$-Minkowski combination given by
$K^{1-t}L^{t}:=\bigcap_{\theta\in S^{n-1}}\left\{x\in{\mathbb{R}}^{n}\colon\langle x,\theta\rangle\leq h_{K}(\theta)^{1-t}h_{L}^{t}(\theta)\right\}$
where $h_{K},h_{L}$ are the support functions of $K$ and $L$ respectively, defined in (19). Owing to the inclusion $K^{1-t}L^{t}\subset(1-t)K+tL$, and the homogeneity of the volume, it is readily clear that (10) implies (1). For more details see Section 4.
Our main contribution to the $\log$-Brunn-Minkowski conjecture is the following equivalence between the inequality (10) and what we call the $\log$-Prékopa-Leindler” inequality.
@@
Let $t\in[0,1]$. Then the Lebesgue measure satisfies inequality (10) for all origin symmetric convex bodies $K$ and $L$, if and only if, for a triple of integrable functions $f,g,h\colon{\mathbb{R}}^{n}\to[0,\infty)$, with $f,g$ even and unimodal functions, such that
@@
(11) $\int_{\mathbb{R}^{n}}h\geq\left(\int_{\mathbb{R}^{n}}f\right)^{1-t}\left(\int_{\mathbb{R}^{n}}g\right)^{t}.$
That is, Theorem 1.3 gives an equivalent functional version of the celebrated $\log$-Brunn-Minkowski conjecture of Böröczky, Lutwak, Yang, and Zhang [10]. In particular, Theorem 1.3 affirmatively answers a question posed by Andrea Colesanti “Is there a functional version of the log-Brunn-Minkowski inequality?” at the conference 61 Probability Encounters, in honour of Sergey Bobkov held at Paul Sabatier University, Toulouse, France in the summer of 2023 [13, Main Question].
Inequality (10) has been verified in several cases. In particular, it has been verified for $n=2$ in the original paper [10]. As consequence of their result and that of Theorem 1.3, the inequality (11) holds when $n=2$ for all even unimodal functions.
In general, our techniques extend inequalities for classes of sets to the class of functions whose sup level sets belong to said sets. For instance in the unconditional case [^3] A function $f\colon{\mathbb{R}}^{n}\to{\mathbb{R}}$ is said to be unconditional if, for each $\varepsilon_{i}\in\{-1,1\}$ $i=1,\dots,n$, and $x=(x_{1},\dots,x_{n})\in{\mathbb{R}}^{n}$, $f(\varepsilon_{1}x_{1},\dots,\varepsilon_{n}x_{n})=f(x)$; a set is unconditional if its characteristic function is an unconditional function., where the log-Brunn-Minkowski conjecture has been verified (see [53]), we have the equivalence [^4] We thank Franck Barthe for the observation that inequality (11) can be deduced from (3) by a well known change of variables (see [53] for more details) in the unconditional case. between the $\log$-Prékopa-Leindler inequality for unconditional functions, inequality (11), and the $\log$-Brunn-Minkowski (10) for $K$ and $L$ unconditional.
As mentioned in the abstract the techniques developed are not limited to Euclidean spaces. In fact, the ideas here are inspired by, and could be used to recover, the proof of a Prékopa-Leindler type inequality on the integers, by the second author and Marsiglietti [47, Theorem 3.8] (see also [29, 28] for more on discrete Prekopa-Leindler-type inequalities). To further hint at broader application, we also give a Borell-Brascamp-Lieb inequality for nilpotent groups, generalizing Tao [56], which in turn generalized the case given for the Heisenberg group [48].
@@
Let $G$ be a simply connected nilpotent Lie group of (topological) dimension $d$ with a Haar measure $\mu$ and $\mathcal{M}_{p}^{(t)}(u,v)$ the $t$-weighted $p$-mean [^5] We refer to the next section and more precisely to (12) for the definition of $\mathcal{M}_{p}^{(t)}$. of $u$ and $v$. Then, for $\alpha\in[-\frac{1}{d},1]$ and $t\in(0,1)$, if
@@
Let us outline the remainder of the article. In Section 2 we introduce an abstract sup-convolution, and derive our main technical tool in Theorem 2.3. In Section 3 we derive the dimensional Brunn-Minkowski inequality for the Gaussian measure and generalizations thereof. In Section 4 we give a functional version of the log-Brunn-Minkowski conjecture and related results for the $L_{p}$-Brunn-Minkowski theory. Finally in Section 5 we discuss further generalizations of the development in Section 2, limitations of our methods, and give a Borell-Brascamp-Lieb type inequality for nilpotent Lie groups.
@@ 2. Generalized Sup-Convolutions
For $u=(u_{1},\dots,u_{n})$ with $u_{i}>0$, $t=(t_{1},\dots,t_{n})$ with $t_{i}>0$ and $p\in(-\infty,0)\cup(0,\infty)$ we define the “mean” ($\sum_{i}t_{i}$ is not necessarily set to be one)
@@
For notational convenience we may sometimes use $\mathcal{M}_{p}^{(t)}(u_{i})$ instead of $\mathcal{M}_{p}^{(t)}(u)$.
@@
is a sup-convolution as well. This is a very particular case of the next result.
@@
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$ and $C(z)=\{x:\Phi(x)=z\}$ for a continuous function $\Phi:E=E_{1}\times\cdots\times E_{n}\to F$. The map defined by
@@
We proceed to verifying superadditivity. Assume that $f=(f_{1},\dots,f_{n})$ and $g=(g_{1},\dots,g_{n})$, with $f_{i},g_{i}$ being functions such that if $g(x_{i})>0$, then $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. Now if $\square g(z)=0$, then the result is trivial from monotonicity. Otherwise, $g(z)>0$, and there exists [^6] Since $g_{i}$ are step functions $x\mapsto\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))$ takes only finite many values. an $x\in C(z)$ such that $\square g(z)=\mathcal{M}_{\alpha}^{(t)}(g_{i}(x_{i}))>0$. For such an $x$, necessarily $g_{i}(x_{i})>0$ so that $f_{i}(x_{i})\geq f_{i}(y_{i})$ for all $y_{i}\in E_{i}$. It follows that
@@
Then $\psi(P)=\Phi(A)$, so that $\Phi(A)$ is analytic, and universally measurable (see [23, Chapter 43]). It follows that $\square s$ is universally measurable as well. ∎
@@
Under the above conventions, given $\mathcal{A}=(\mathcal{A}_{1},\dots,\mathcal{A}_{n})$ with each $\mathcal{A}_{i}$ a sub-collection of analytic sets, for $\alpha\leq 1$ and $t=(t_{1},\dots,t_{n})$, with $t_{i}>0$ such that $\sum_{i=1}^{n}t_{i}=1$, a triple $(\nu,\mu,\square)$ and $p=(p_{1},\dots,p_{n})\in[1,\infty)^{n}$ satisfies
(14) $\nu(\square a)\geq\mathcal{M}_{\alpha}^{(t)}(\|a_{i}\|_{p_{i}})$
@@
(15) $\nu_{*}(\square f)\geq\mathcal{M}_{\alpha}^{(t)}(\|f_{i}\|_{p_{i}}),$
@@
Before we prove the theorem, we require the following lemma.
@@
Assume that $p_{i}=1$, for each $i=1,\ldots,n$. Let us first observe that it is enough to prove the inequality for $f_{i}$ step functions. Define
@@
$\nu_{*}\left(\square f\right)\geq\nu\left(\square f^{(m)}\right)\geq\mathcal{M}_{\alpha}^{(t)}\left(\|f^{(m)}\|_{1}\right),$
@@
$\mathcal{M}_{\alpha}^{(t)}(\|f\|_{1})=\lim_{m\to\infty}\mathcal{M}_{\alpha}^{(t)}(\|f^{(m)}\|_{1})\leq\nu(\square f).$
@@
Note that $f_{1}^{-}$ and $f_{1}^{+}$ individually take strictly fewer values than $f_{1}=f_{1}^{-}+f_{1}^{+}$. Define
$\lambda=\frac{\|f^{-}_{1}\|_{1}}{\|f_{1}\|_{1}},$
so that $1-\lambda=\frac{\|f^{+}_{1}\|_{1}}{\|f_{1}\|_{1}}$. Note that $\lambda\in(0,1)$. Let
$f_{i}^{-}=\min\{f_{i},s_{i}\}\ \hbox{ and }\ f_{i}^{+}=\max\{f_{i}-s_{i},0\}$
where $s_{i}$ is implicitly and uniquely defined through the equation
$\frac{\|f_{i}^{-}\|_{1}}{\|f_{i}\|_{1}}=\lambda$
because $s\mapsto\mu_{i}(\min\{f_{i},s\})$ is continuous and strictly increasing on $(0,\|f_{i}\|_{\infty})$. Note that $f_{i}=f_{i}^{-}+f_{i}^{+}$ and that $1-\lambda=\frac{\|f_{i}^{+}\|_{1}}{\|f_{i}\|_{1}}$, $i=1,2,\dots,n$.